OpenStudy (anonymous):

Find the sum of the first 12 terms of the sequence. 1, -4, -9, -14, . . .

OpenStudy (yoloshroom):

alright so, first we have to figure out the sequence, what do you think it is so far? Hint : 1 + x = -4 , -4 + x = -9 ....

OpenStudy (anonymous):

-9+x=-14?

OpenStudy (anonymous):

so -5

OpenStudy (yoloshroom):

Yup! so now that we know how to fnd the next numbers in the sequence we can proceed to find all 12 terms, and add them up

OpenStudy (yoloshroom):

What did you get? :)

OpenStudy (anonymous):

is it -338?

OpenStudy (yoloshroom):

1+(-4)-9-14-19-24-29-34-39-44-49-54 = = -319 i believe. Let me double check my work :)

OpenStudy (yoloshroom):

I believe it's -319, mind showing me what you did?

OpenStudy (anonymous):

1,-4,-9,-14,-19,-24,-29,-34,-39,-44,-59,-64 , my bad i accidently typed in this i didnt do 54

OpenStudy (yoloshroom):

alright, but you understand it right? :) cuz if u need clarification i can easily help you put with that